OpenStudy (anonymous):

Find the slope and y-intercept of the line in the given equation. y=-1/5x-6

OpenStudy (anonymous):

m=-1/5 y-intercept=-6

OpenStudy (anonymous):

y-intercept whent x=0, y=-6 and y=mx+c so m=-1/5

OpenStudy (anonymous):

since it's in the form y=mx+b and m=slope and b=y-intercept, slope=-1/5 and y-intercept=-6